# Optimal Control of the Future via Prospective Foraging

### v1

@@
Optimal Control of the Future via Prospective Foraging
@@
Optimal control of the future is the next frontier for AI. Current approaches to this problem are typically rooted in either reinforcement learning or online learning. While powerful, these frameworks for learning are mathematically distinct from Probably Approximately Correct (PAC) learning, which has been the workhorse for the recent technological achievements in AI. We therefore build on the prior work of prospective learning, an extension of PAC learning (without control) in non-stationary environments [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik]. Here, we further extend the PAC learning framework to address learning and control in non-stationary environments. Using this framework, called “Prospective Control”, we prove that under certain fairly general assumptions, empirical risk minimization (ERM) asymptotically achieves the Bayes optimal policy. We then consider a specific instance of prospective control, foraging—which is a canonical task for any mobile agent—be it natural or artificial. We illustrate that existing reinforcement learning algorithms fail to learn in these non-stationary environments, and even with modifications, they are orders of magnitude less efficient than our prospective foraging agents. Code is available at: https://github.com/neurodata/ProspectiveLearningwithControl.
@@ Introduction
When building real-world machine learning and/or artificial intelligence (ML/AI) systems, we almost always care about their future performance because we will be using them in the future. And yet, many classical formulations of ML/AI essentially ignore this fact. Consider, for example, Probably Approximately Correct (PAC) learning [Mohri2012-qj], the primary workhorse of modern ML/AI. The key assumption underlying this framework is that each data point is sampled independently and identically from a given fixed distribution (IID). In other words, the future is assumed to be identical to the past (up to distribution) and also independent of any past data. This assumption is woefully inadequate for many real-world problems.
Reinforcement learning (RL) and online learning (OL) are two popular machine learning paradigms relaxing this IID assumption. Rather than IID, RL typically models the environment as a Markov Decision Process (MDP) with states, actions, rewards, and aims to learn a behavior/policy that maximizes long-term return [Bertsekas2023, Sutton2018-gx, StrehlLiLittman2009PACMDPs]. RL has achieved many real-world successes, e.g. Go [Silver2016NatureGo]. However, standard RL methods fail in non-stationary reward settings and require multiple episodes (though see [Chen2022-ja]). On the other hand, online learning algorithms are run in a sequence of consecutive rounds, making a prediction, observing feedback, and updating [ShalevShwartz2012]. The OL framework tends to make no distributional assumptions about the data whatsoever. This is a worst-case assumption scenario, because it essentially says it knows absolutely nothing about the future, the opposite of IID. Thus, the goal of online learning algorithms is to find the best fixed hypothesis in hindsight.
Neither of these frameworks, therefore, relaxes the IID assumption in a fully satisfactory way. Prospective Learning (PL) [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik] introduced recently generalizes PAC learning to address stochastic processes, rather than independent and identically distributed data. In those prior works, the IID assumption was relaxed, but decisions had no impact on the future. Here, we generalize further to incorporate stochastic decision processes, so that future loss and data distributions also depend on current actions.
As a concrete example, consider foraging. Foraging is informally defined as the sequential search for resources [Barack2024-sx]. It is a central behavior of all mobile agents, be they living organisms, AI bots, or robots. Standard formalizations of foraging have significant limitations in terms of characterizing real-world behaviors. Most importantly for this work, classical models of foraging assume that the environment is essentially stationary. For example, without the single agent’s intervention, resources stay available indefinitely. This is a poor model of real-world foraging because, typically, resources are depleted, rather than persistent.
We therefore introduce prospective learning with control (PL+C), as a generalization to the prior work on prospective learning without control (PL-C) [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik]. We first briefly summarize PL-C using a slightly revised notation (2). Then we formally introduce PL+C (3). Our framework enables us to prove the existence of a sequence of hypotheses whose risks converge to the Bayes optimal risk (4). We then introduce our Prospective Foraging (ProForg) (5). ProForg performs substantially better than classical RL algorithms—even with suitable modifications to address the non-stationarity—on a dynamic foraging problem (7). We therefore believe that PL+C presents a promising strategy for learning to control the future optimally and efficiently.
Prospective Learning without control (PL-C)
If $z$ is a discrete set, the above integral becomes a sum. Let $p_{z}$ denote the probability of a particular sequence $z\in\mathbb{Z}$. With this notation, Equation˜1 becomes
Empirical Goal In practice, in general, we cannot minimize Equation˜1 because it requires solving an intractable integral, and we typically cannot solve Equation˜2, because it requires sampling too many possible futures (there are exponentially many even if $z$ is binary). Moreover, we cannot minimize Equations˜1 and 2 because they both require sampling from an unknown distribution, $\mathbb{P}_{Z\mid z_{\leq t}}$. To address the first issue, we use a Monte Carlo approximate sum, sampling only a subset of possible futures. To address the second issue, rather than sampling from the unobserved future, we sample only from the observed past. With these two approximations in hand, we have
Main result The main result of De-Silva2024-jj is approximating a variant of Equation˜1, under suitable assumptions, satisfies our theoretical goal. The key assumptions are (a) consistency and (b) uniform concentration. The focus in that work was on settings in which our decisions $\hat{h}^{t}(x_{s},s)$ had no impact on future outcomes or distributions. In such cases, we can ignore the sum in Equation˜1, because all that matters for the current decision is the instantaneous loss. Here, we are interested in general settings where the decisions also impact future distributions and losses.
Prospective Learning with Control (PL+C)
We modify the above prospective learning framework to accommodate stochastic decision processes, rather than merely stochastic processes. In other words, we extend the framework from learning to predict the future, to learning to control the future. The key difference is in the loss. For prediction, the loss is the error relative to the truth/optimal, which is always observed. In control, we do not necessarily see the truth/optimal, so one cannot define loss as an error term. Below, we formalize this distinction.
@@
Data Let $x_{t}$ be the position of the agent in an environment at time $t$. Thus, if the environment is two-dimensional, finite, rectangular, and discrete, then it has width $W$ and height $H$, and $x_{t}\in\mathcal{X}\equiv[W]\times[H]$ is the state of the agent at time $t$. As above, we also record time $t\in\mathcal{T}\subset\mathbb{N}$. We refer to the position and time together as the observation state (or simply state), $(x_{t},t)$. Then $y_{t}\in\mathcal{Y}$ is the reward the agent would receive at time $t$ at any of the positions. Thus, if position is a two-dimensional grid as described above, then $\mathcal{Y}\equiv\mathbb{R}_{+}^{W\times H}$. Thus, the past data are the observation state (position and time) and observed rewards.
Hypothesis The hypothesis maps from the current state to the next position, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}$. Without an estimate of the rewards, however, we have no way, in general, to estimate what the rewards would be if we changed the hypothesis. Thus, we could either let the hypothesis map to the next position and the next reward, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}\times\mathbb{R}_{+}$, and so $h(x_{s},s)\to x_{s+1},y_{s+1}(x_{s+1})$. More generally, the hypothesis maps to the next position and the reward at all possible positions, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}\times\mathcal{Y}$, so $h(x_{s},s)\to(x_{s+1},y_{s+1})$.
@@
Loss The instantaneous loss operates on the current hypothesis and data, $\bar{\ell}:\mathcal{H}\times\mathbb{Z}\to\mathbb{R}_{+}$. The instantaneous loss is the negative reward the agent receives at the position the agent is in at that time. Recalling that $x_{s}$ encodes the position of the agent at time $s$, and that $y_{s+1}$ encodes the reward at all locations at time $s+1$, so, in a slight abuse of notation, we can write
@@
The loss at time $t$ is the cumulative instantaneous loss, $\sum_{s>t}w_{s-t}\bar{\ell}(h_{s}(x_{s},s),y_{s+1})$. [^7] We intentionally do not specify whether the horizon is finite or infinite here.
@@
Assumptions Let $\bm{X}=X_{0,1,2,\dots}$ and $\bm{Y}=Y_{0,1,2,\dots}$ be stochastic decision processes, that is, $(\bm{X},\bm{Y})\sim F\in\mathcal{F}$, where $Y_{>t}$ can be an arbitrary function of $(X_{\leq t},Y_{\leq t})$. Denote the values of the random variables $\bm{X}$ and $\bm{Y}$ as $\bm{x}$ and $\bm{y}$, respectively. Note that the distribution of $X_{t}$ is, in general, a function of the previous sequence of observed $x_{\leq t}$, which is itself a function of the observed rewards $y_{\leq t}$, and the estimated hypotheses, $\hat{h}^{\leq t}$. If we assume that the learners and hypotheses are deterministic functions of the data, then future distributions of $X_{s}$ are functions of the learner and hypotheses, but not conditionally independent on them, once conditioned on the data.
Theoretical Goal The goal of learning a hypothesis whose risk converges to the Bayes optimal risk remains the same in prospective learning with or without control. However, the objective function we seek to optimize is slightly modified because control renders the problem inherently sequential. Thus, the instantaneous loss in PL without control (Equation˜1) is replaced with the instantaneous loss in PL with control (Equation˜4):
$\hat{h}^{t}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}R_{t}(h)\equiv\operatornamewithlimits{argmax}_{h\in\mathcal{H}_{t}}\int\sum_{s>t}w_{s-t}\times y_{s+1}(x_{s+1})d\mathbb{P}_{Z\mid z_{\leq t}}=\operatornamewithlimits{argmax}_{h\in\mathcal{H}_{t}}\sum_{z\in\mathbb{Z}:z_{\leq t}}p_{z}\sum_{s>t}w_{s-t}\times y_{s+1}(x_{s+1}).$ (5)
Note that while the $\operatornamewithlimits{argmax}$ is with respect to $h$, on the right-hand side after the $\equiv"$, the $h$ is implicit. We could notationally let $y_{s}$ be a function of $h$ to clarify if we desired.
Empirical Goal Much like PL without control, in PL with control, we cannot directly solve Equation˜5 because we cannot sample all possible futures, and we do not know the likelihood of each possible future. Thus, we modify the sums in Equation˜5 to only sum up to the current time, $t$:
@@
Relationship to prospective learning without control The observed data triples in PL+C are $(x_{s},y_{s+1}(x_{s+1}),s)$, rather than $(x_{s},y_{s},s)$ in PL-C. Note that these data triples differ in two ways. First, the index on $y$ changes from $s$ to $s+1$ as we go from PL-C to PL+C; this is required because the rewards are necessarily a function of the action of the agent. Second, rather than the entire vector of output being observed as in PL-C, in PL+C, we only obtain a partial observation: the reward at the location we happen to currently inhabit. This renders PL+C closely related to causal inference, in which the outcomes are defined as a function of the intervention (or treatment), and therefore only partially observed, leading to the potential outcomes framework [Neyman1923-aw, Rubin1974-gt]. Hypotheses are no longer maps from $\mathcal{X}\times\mathcal{T}$ to $\mathcal{Y}$, rather, they map to $\mathcal{X}$ and $\mathbb{R}_{+}$. Loss in PL with or without control is the cumulative instantaneous loss. However, in PL-C, instantaneous loss is typically an error signal, e.g., the difference between the predicted output and the actual output. In PL+C, loss is just a negative reward. Moreover, loss is not observed for any of the decisions that the agent did not make, necessitating the use of some surrogate loss function in practice. Learners map from data to hypotheses, though the data are subtly different due to the time indices and partial observation. Assumptions can be the same, though note that the future data are now necessarily a function of the learner and hypothesis in PL+C, whereas they were not necessarily in PL-C. The theoretical goal is the same, modulo replacing the loss from PL-C with that from PL+C. The empirical goal changes somewhat similarly, insofar as in PL+C we cannot directly solve Equation˜5, so we use a surrogate optimization problem, Equation˜6.
Theoretical Results
Here we provide proof sketches; formal proofs are in Appendix˜A.
Lemma 4.1.
Theorem 4.1.
Then, there exists a sequence $\{i_{t}\}_{t=1}^{\infty}$ such that
Estimation & inference for prospective learning with control
Warm Start via Batch Pretraining
Cumulative loss estimation
Recall the empirical goal in prospective learning with control, Equation˜6. Based on this equation, we build a data corpus using prospective loss as follows. Let $\widetilde{x}_{s}\in\mathbb{R}^{d_{p}}$ and $\widetilde{s}\in\mathbb{R}^{d_{t}}$ be the embedded positions and times, and let $d=d_{p}+d_{t}$. Let $X_{t}=(\widetilde{x}_{s},\widetilde{s})_{s\leq t}$ be the input matrix, such that $X_{t}\in\mathbb{R}^{d\times t}$. Let $Y_{t}\in\mathbb{R}^{t}$ be the sequence of cumulative rewards, that is $Y_{t}=(\sum_{t^{\prime}=s+1}^{t}-w_{s-t^{\prime}}\times y_{t}(x_{t^{\prime}}))_{s\leq t}$. Then, we can feed data into any supervised learning machine algorithm, for example, a decision forest or a deep network. Let the resulting regression function be denoted $\hat{g}^{p}$, which maps from a state, $(x_{s},s)$ to an estimated prospective loss $\bar{y}_{s+1}\in\mathbb{R}_{+}$.  [^8] We implicitly assume that $\hat{g}$ implements the embedding function.
Instantaneous loss estimation
When searching over hypotheses, some will result in positions at various times that are not identical (i.e., counterfactual) to the positions that were selected by the estimated hypothesis at that time, $\hat{h}^{s}$. We therefore lack observations of $y_{s}$ for all the counterfactual positions, and in practice, we cannot actually solve Equation˜6, because we have missing data. Instead, we introduce a further approximation, replacing $y_{s+1}$ with $\tilde{y}_{s+1}$. Define $\widetilde{y}_{s}$ as follows. Let $x^{\prime}_{s}$ denote potential outcomes that the agent can take at time $s$, whereas $x_{s}$ is the position that the agent actually took at that time, and let $\hat{y}_{s}(x^{\prime}_{s})$ denote the estimated reward at time $s$ at position $x^{\prime}_{s}$ based on the current hypothesis:
Online estimation and inference
Recall that our goal is Equation˜6. And now, we get estimated instantaneous reward $\tilde{y}_{s}$, and cumulative reward $\bar{y}_{s}$, respectively. We combine these two parts to approximate the infinite horizon prospective loss. This yields both counterfactual insight and a tractable approximation to future loss. To be more specific, given a predefined horizon, we have $\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\{w_{t+T-t^{\prime}}\times\bar{y}_{t+T}+\sum_{s>t^{\prime}}^{t+T}-w_{s-t^{\prime}}\times\tilde{y}_{s+1}(x_{s+1})\}$, where $\bar{y}_{t+T}$ represents an estimation of terminal cost [Bertsekas2023] at time $t+T$, and $\tilde{y}_{s}$ is the estimated instantaneous cost at time $s$. Here, two supervised learning models are used: one for approximating the look-ahead cost for the finite steps, and another to approximate the terminal cost correcting the truncation of large or infinite horizons.
Inference in Practice
We evaluate each potential future trajectory via calculating its prospective loss, $\{w_{T}\times\hat{g}^{p}(x_{s+T},s+T)+\sum_{j=s}^{s+T}w_{j-s}\times\hat{g}^{i}(x_{s},s)\}$. Assuming $T=6$, and $A=2$, we have $2^{6}=64$ different trajectories, each with an associated prospective loss, as estimated by iterating the instantaneous losses and adding the terminal cost at the end. We select the next state that has the trajectory with the maximum value. For each time $t$, the decision is made by the learner trained on past observations till time t1t-1. Algorithm 1 in Appendix 6 provides details. The above is an exhaustive search strategy. If our computational budget is inadequate to search all the possible trajectories, then one can search a subset of possible trajectories using Monte Carlo Tree Search (MCTS) [Bruegmann1993mcgo, Kocsis2006uct, Coulom2007mcts, Silver2016NatureGo, silver2017masteringchessshogiselfplay].
Algorithm
@@
Experimental Validation
To validate the above-described theoretical framework, we devise a specific prospective learning with control problem, namely, prospective foraging.
@@
We implement and evaluate the prospective forager (ProForg) within the 1-D foraging environment described in Section˜7.1 and Bai2026-ik: the agent forages along a $1\times 7$ linear track containing only two reward patches, A and B, located three grid cells apart, as shown in Figure˜1. Every $r=10$ timestamps, rewards alternatively appear at A and B. The agent’s goal is to maximize the total reward in its single lifetime, meaning there are no resets within each run.
We define the prospective regret of a hypothesis $h^{(t)}$ as the difference between the prospective risk returned by hypothesis $h^{(t)}$ and Bayes optimal risk: $\Delta{h^{(t)}}=R_{t}(h^{(t)})-R_{t}^{*}$. In practice, one cannot evaluate regret over an infinite future. Therefore, we approximate the prospective regret by evaluating it over a finite horizon of $T$ time steps and returning the normalized one (NPR): $\frac{1}{T}\Big(R_{t,T}\!\big(h^{(t)}\big)-R_{t,T}^{\star}\Big)$, where $R_{t,T}$ is the finite horizon prospective risk. In addition, since we only have one trajectory after time t, we replace the risk by the prospective loss, so $\Delta{h^{(t)}}\approx\frac{1}{T}\sum_{i=s}^{s+T}w_{i-s}(y_{s}(x^{*}_{s})-y_{s}(x_{s}))$, where $(x^{*}_{s},x^{*}_{s+1},....,x^{*}_{s+T})$ are the states returned by Bayes optimal hypothesis at the corresponding times.
@@
After an initial warm-up phase, we let the ProForg learn in an online manner as it interacts with the environment over time (Figure˜2). The procedure follows the steps detailed in Algorithm˜1. During the online updating part, we retrained regressors with all data collected so far for each update. We compute the normalized prospective regret of ProForg over time, and compare it to Fitted Q-Iteration (FQI) [ernst2005tree, munos2008finite] and Soft Actor-Critic (SAC) [haarnoja2018soft, haarnoja2018softa], two classical reinforcement learning algorithms, and their time-aware variants that we introduce here to enable them to work in non-stationary environments (see Appendices˜B, 2 and 3). In both ProForg and FQI variants, we use Random Forests as our chosen class of regressors.
@@
Discussion

### v2

# Optimal control of the future via prospective learning with control

@@
Optimal control of the future via
prospective learning with control
@@
Optimal control of the future is the next frontier for AI. Current approaches to this problem are typically rooted in either reinforcement learning (RL). While powerful, this learning framework is mathematically distinct from supervised learning, which has been the main workhorse for the recent achievements in AI. Moreover, RL typically operates in a stationary environment with episodic resets, limiting its utility to more realistic settings. Here, we extend supervised learning to address learning to control in non-stationary, reset-free environments. Using this framework, called ”Prospective Learning with Control (PL+C)”, we prove that under certain fairly general assumptions, empirical risk minimization (ERM) asymptotically achieves the Bayes optimal policy. We then consider a specific instance of prospective learning with control, foraging—which is a canonical task for any mobile agent—be it natural or artificial. We illustrate that modern RL algorithms fail to learn in these non-stationary reset-free environments, and even with modifications, they are orders of magnitude less efficient than our prospective foraging agents.
@@ Introduction
When building artificial intelligence (AI) systems, we often start with some basic operating assumptions about the world. For example, a prevailing, though obviously wildly inaccurate assumption, is that time does not exist. This is a core assumption in much of foundational estimation and learning theory dating back nearly a century [Glivenko1933-wm, Cantelli1933-vj, Vapnik1971-um, Valiant1984-dx], all the way through modern machine learning tools including large language models [vaswani2017attention]. A second prevailing and obviously wildly inaccurate assumption is that decisions have no impact. This assumption leads to developing systems based purely on predictions and forecasts, without concerning oneself with the consequences of one’s actions. For example, consider early machine learning successes such as spam detection [Sahami1998-yr], which operated according to both assumptions: no time and no impact. Once it is deployed, spammers change their behavior to subvert the system. This simple example illustrates a few key properties required of deployed machine learning systems.
First, time is real. In particular, the world is dynamic; therefore, for the system to preserve its desirable properties, it must also be dynamic on approximately the same time-scales. Second, decisions have impacts. Therefore, pure inference or forecasting systems are missing a key ingredient. The history of machine learning has traversed a trajectory of starting with the simplest possible assumptions, and slowly relaxing these assumptions to obtain more effective deployed systems. One limitation of relaxing the simplifying assumptions is that as we strive to estimate more complex functions, we require more data. Thus, as our ability to acquire and analyze more high-quality data has improved, so has our toolbox generalized accordingly.
Nearly one hundred years ago, people realized that one could estimate arbitrary properties of data under quite general assumptions, as long as we assumed no time and no impact [Glivenko1933-wm, Cantelli1933-vj]. In the 1950s, decision theory seriously relaxed the assumption of no impact [Wald1949-fl]. Markov models and Hidden Markov Models (HMM) relaxed the assumption of no time [Baum1966-ep], and Markov decision processes (MDP) and Partially Observable MDPs (POMDPs) further relaxed the assumption of no impact with the Bellman equation [Bellman1958-us], introducing control theory. Then Kalman filters addressed the fact that time and impact were real, but the parameters of the system were partially unknown [Kalman1960-mm], founding adaptive control [Landau1984-mr]. Modern approaches to adaptive control are typically considered reinforcement learning (RL) [Sutton2018-gx]. Modern ML has many successes, though they are typically restricted to settings where the environment is stationary, there are many trials, and the whole world is perfectly observable (e.g., in video games [Silver2016NatureGo]). Non-stationary adaptive control and RL, and continual RL address the possibility of an environment that changes over time [Khetarpal2022-tc, Abel2023-gl, Kumar2025-go]. However, they typically operate under the assumption of many trials, rather than single-life [Chen2022-dq] and reset-free [Lu2020-bp] approaches. As is the norm in RL, these approaches continue to operate under a Markov Decision Process framework.
Our contribution builds on these earlier works, addressing non-stationary reset-free environments without making an MDP assumption. In a departure from much of RL, our work builds on estimation theory [Glivenko1933-wm, Cantelli1933-vj], Probably Approximately Correct learning theory [Valiant1984-dx, Valiant2013-gp], and modern machine learning and supervised learning practice [Mohri2012-qj, Goodfellow2016-kf]. Continual learning [Vogelstein2025-uq] and prospective learning [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik] generalize estimation and learning theory by relaxing the assumption of no time. In this work, we introduce Prospective Learning with Control (PL+C), in which we further relax the assumption of no impact.
We formally introduce our framework in Section˜2. Algorithms for solving PL+C problems are provided in Appendix˜B. Numerical results are provided in Section˜4, demonstrating that leading RL algorithms, even augmented to incorporate non-stationarity, either fail to converge to the Bayes optimal, or converge orders of magnitude more slowly than our algorithms. Theory proving that our algorithms converge to the Bayes optimal solution under suitable assumptions is in Section˜5. This collection of empirical and theoretical results is compelling evidence that prospective learning with control is worthy of further study.
Framework for Prospective Learning with Control (PL+C)
This work builds on prospective learning defined previously [De_Silva2023-al, De-Silva2024-jj, Bai2026-ik], which we refer to hereafter as Prospective Learning without Control (PL-C). See Appendix˜A for a review of that work with slightly modified notation, and contrast with Prospective Learning with Control (PL+C), as defined below.
@@
Data Let $x_{t}$ be the position of the agent in an environment at time $t$. We also record time $t\in\mathcal{T}\subset\mathbb{N}\equiv[1,2,3,\dots]$. We refer to the position and time together as the observation state (or simply state), $(x_{t},t)$. Let $y_{t}\in\mathcal{Y}$ be the reward the agent would receive at time $t$ at any of the positions.
Hypothesis The hypothesis maps from the current state to the next position, $h:\mathcal{X}\times\mathcal{T}\to\mathcal{X}$.
@@
Loss The instantaneous loss is the negative reward the agent receives at the position the agent is in at that time. Recalling that $x_{s}$ encodes the position of the agent at time $s$, and that $y_{s+1}$ encodes the reward at all locations at time $s+1$, we can write
@@
The loss at time $t$ is the weighted cumulative instantaneous loss, $\sum_{s>t}w_{s-t}\bar{\ell}(h_{s}(x_{s},s),y_{s+1})$. [^1] We intentionally do not specify whether the horizon is finite or infinite here.
@@
Theoretical Goal Define risk as the expected loss, $R_{t}(h)=\operatorname*{\mathbb{E}}_{F}[\ell_{t}(h,Z)\mid z_{\leq t}]=\int\ell_{t}(h,Z)d\mathbb{P}_{Z\mid z_{\leq t}}$. For risk to be well defined, we assume that $\bm{x}$ and $\bm{y}$ are realizations of random variables, $\bm{X}$ and $\bm{Y}$, which are jointly distributed as $(\bm{X},\bm{Y})\sim F\in\mathcal{F}$.  [^2] Note that the distribution of $\bm{Y}_{>t}$ and $\bm{X}_{>t}$ are, in general, functions of the previous sequence of observed $x_{\leq t}$, which is itself a function of the observed rewards $y_{\leq t}$, and the estimated hypotheses, $\hat{h}^{\leq t}$. The goal is to choose a learner that obtains a hypothesis that minimizes risk.
$\hat{h}^{t}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}_{t}}R_{t}(h)\equiv\operatornamewithlimits{argmax}_{h\in\mathcal{H}_{t}}\int\sum_{s>t}w_{s-t}\times y_{s+1}^{h}(x_{s+1}^{h})d\mathbb{P}_{Z\mid z_{\leq t}},$ (2)
where the notation $y^{h}_{s+1}$ and $x^{h}_{s+1}$ indicate that the subsequent positions and rewards are a function of $h$.
Empirical Goal We cannot directly solve Equation˜2 because we cannot sample all possible futures, and we do not know the likelihood of each possible future. We therefore replace the integral in Equation˜2 with a sum up to the current time, $t$:
@@
Algorithms for Prospective Learning with Control
@@
Warm Start via Batch Pretraining
Cumulative loss estimation
Recall the empirical goal in prospective learning with control, Equation˜3. Based on this equation, we build a data corpus using prospective loss as follows. Let $\widetilde{x}_{s}\in\mathbb{R}^{d_{p}}$ and $\widetilde{s}\in\mathbb{R}^{d_{t}}$ be the embedded positions and times, and let $d=d_{p}+d_{t}$. Let $X_{t}=(\widetilde{x}_{s},\widetilde{s})_{s\leq t}$ be the input matrix, such that $X_{t}\in\mathbb{R}^{d\times t}$. Let $Y_{t}\in\mathbb{R}^{t}$ be the sequence of cumulative rewards, that is $Y_{t}=(\sum_{t^{\prime}=s+1}^{t}-w_{s-t^{\prime}}\times y_{t}(x_{t^{\prime}}))_{s\leq t}$. Then, we can feed data into any supervised learning machine algorithm, for example, a decision forest or a deep network. Let the resulting regression function be denoted $\hat{g}^{p}$, which maps from a state, $(x_{s},s)$ to an estimated prospective loss $\bar{y}_{s+1}\in\mathbb{R}_{+}$.  [^3] We implicitly assume that $\hat{g}$ implements the embedding function.
Instantaneous loss estimation
When searching over hypotheses, some will result in positions at various times that are not identical (i.e., counterfactual) to the positions that were selected by the estimated hypothesis at that time, $\hat{h}^{s}$. We therefore lack observations of $y_{s}$ for all the counterfactual positions, and in practice, we cannot actually solve Equation˜3, because we have missing data. Instead, we introduce a further approximation, replacing $y_{s+1}$ with $\tilde{y}_{s+1}$. Define $\widetilde{y}_{s}$ as follows. Let $x^{\prime}_{s}$ denote potential outcomes that the agent can take at time $s$, whereas $x_{s}$ is the position that the agent actually took at that time, and let $\hat{y}_{s}(x^{\prime}_{s})$ denote the estimated reward at time $s$ at position $x^{\prime}_{s}$ based on the current hypothesis:
Online estimation and inference
Recall that our goal is Equation˜3. And now, we get estimated instantaneous reward $\tilde{y}_{s}$, and cumulative reward $\bar{y}_{s}$, respectively. We combine these two parts to approximate the infinite horizon prospective loss. This yields both counterfactual insight and a tractable approximation to future loss. To be more specific, given a predefined horizon, we have $\hat{h}^{t}\approx\operatornamewithlimits{argmin}_{h^{t}\in\mathcal{H}_{t}}\sum_{t^{\prime}>0}^{t}\{w_{t+T-t^{\prime}}\times\bar{y}_{t+T}+\sum_{s>t^{\prime}}^{t+T}-w_{s-t^{\prime}}\times\tilde{y}_{s+1}(x_{s+1})\}$, where $\bar{y}_{t+T}$ represents an estimation of terminal cost [Bertsekas2023] at time $t+T$, and $\tilde{y}_{s}$ is the estimated instantaneous cost at time $s$. Here, two supervised learning models are used: one for approximating the look-ahead cost for the finite steps, and another to approximate the terminal cost correcting the truncation of large or infinite horizons.
Inference in Practice
We evaluate each potential future trajectory via calculating its prospective loss, $\{w_{T}\times\hat{g}^{p}(x_{s+T},s+T)+\sum_{j=s}^{s+T}w_{j-s}\times\hat{g}^{i}(x_{s},s)\}$. Assuming $T=6$, and $A=2$, we have $2^{6}=64$ different trajectories, each with an associated prospective loss, as estimated by iterating the instantaneous losses and adding the terminal cost at the end. We select the next state that has the trajectory with the maximum value. For each time $t$, the decision is made by the learner trained on past observations till time t1t-1. Algorithm 1 provides details. The above is an exhaustive search strategy. If our computational budget is inadequate to search all the possible trajectories, then one can search a subset of possible trajectories using Monte Carlo Tree Search (MCTS) [Bruegmann1993mcgo, Kocsis2006uct, Coulom2007mcts, Silver2016NatureGo, silver2017masteringchessshogiselfplay].
Experiments for Prospective Learning with Control
To validate the above-described theoretical framework, we devise a specific prospective learning with control problem, namely, prospective foraging. Foraging is informally defined as the sequential search for resources [Barack2024-sx]. It is a central behavior of all agents, be they living organisms, AI bots, or robots. Standard formalizations of foraging assume that the environment is essentially stationary. For example, without the single agent’s intervention, resources stay available indefinitely. This is a poor model of real-world foraging because, typically, resources are depleted, rather than persistent. We therefore introduce prospective foraging, as a special instance of a prospective learning with control problem.
@@
We implement and evaluate the prospective forager (ProForg) within the 1-D foraging environment described in Section˜4.1 and Bai2026-ik: the agent forages along a $1\times 7$ linear track containing only two reward patches, A and B, located three grid cells apart, as shown in Figure˜1. Every $r=10$ timestamps, rewards alternatively appear at A and B. The agent’s goal is to maximize the total reward in its single lifetime, meaning there are no resets within each run.
We define the prospective regret of a hypothesis $h^{(t)}$ as the difference between the prospective risk returned by hypothesis $h^{(t)}$ and Bayes optimal risk: $\Delta{h^{(t)}}=R_{t}(h^{(t)})-R_{t}^{*}$. In practice, one cannot evaluate regret over an infinite future. Therefore, we approximate the prospective regret by evaluating it over a finite horizon of $T$ time steps and returning the average one (Averaged Prospective Regret): $\frac{1}{T}\Big(R_{t,T}\!\big(h^{(t)}\big)-R_{t,T}^{\star}\Big)$, where $R_{t,T}$ is the finite horizon prospective risk. In addition, since we only have one trajectory after time t, we replace the risk by the prospective loss, so $\Delta{h^{(t)}}\approx\frac{1}{T}\sum_{i=s}^{s+T}w_{i-s}(y_{s}(x^{*}_{s})-y_{s}(x_{s}))$, where $(x^{*}_{s},x^{*}_{s+1},....,x^{*}_{s+T})$ are the states returned by Bayes optimal hypothesis at the corresponding times.
@@
After an initial warm-up phase, we let the ProForg learn in an online manner as it interacts with the environment over time (Figure˜2). The procedure follows the steps detailed in Algorithm˜1. During the online updating part, we retrained regressors with all data collected so far for each update. We compute the normalized prospective regret of ProForg over time, and compare it to Fitted Q-Iteration (FQI) [ernst2005tree, munos2008finite] and Soft Actor-Critic (SAC) [haarnoja2018soft, haarnoja2018softa], two classical reinforcement learning algorithms, and their time-aware variants that we introduce here to enable them to work in non-stationary environments (see Appendices˜C, 2 and 3). In both ProForg and FQI variants, we use Random Forests as our chosen class of regressors.
@@
Theory for Prospective Learning with Control
Here we provide proof sketches; formal proofs are in Appendix˜B.
Lemma 5.1.
Theorem 5.1.
Then, there exists a sequence $\{i_{t}\}_{t=1}^{\infty}$ such that
Discussion of Prospective Learning with Control